TWO ANGLES ARE Complementary the sum of measure of the first angle and half of the second is 69 find the measures of the angles
what is the measure of the smaller angle? what is the measure of the other angle?

Answer by mananth(16949)

let the angle be x
the other angle will be 90-x
..
x+1/2 *(90-x)=69
LCM =2
2x+90-x /2 = 69
Multiply by2
x+90=2*69
x+90=138
add -90
x+90-90=138-90
x=48 deg one angle
the other angle will ne 90 -48=42 deg